# Accepted at AEQUITAS 2024: Workshop on Fairness and Bias in AI | co-located with ECAI 2023, Santiago de Compostela, Spain

### v3

@@ Abstract
We present FairX, an open-source Python-based benchmarking tool designed for the comprehensive analysis of models under the umbrella of fairness, utility, and eXplainability (XAI). FairX enables users to train benchmarking bias-removal models and evaluate their fairness using a wide array of fairness metrics, data utility metrics, and generate explanations for model predictions, all within a unified framework. Existing benchmarking tools do not have the way to evaluate synthetic data generated from fair generative models, also they do not have the support for training fair generative models either. In FairX, we add fair generative models in the collection of our fair-model library (pre-processing, in-processing, post-processing) and evaluation metrics for evaluating the quality of synthetic fair data. This version of FairX supports both tabular and image datasets. It also allows users to provide their own custom datasets. The open-source FairX benchmarking package is publicly available at https://github.com/fahim-sikder/FairX.
@@ 1 Introduction
With the rapid development of artificial intelligence-based systems to aid us in our daily lives, it is important for these systems to give outcomes that is acceptable for all users, including—but not limited to—from demographic perspective. Troublingly, as the available data is filled with human or machine bias, models trained with these dataset often gives unfair outcome towards some demographic [1]. It is therefore critical to mitigate bias in the dataset and model. Over the years, researchers have used different techniques to achieve this. These techniques can be roughly grouped in three families: 1) Pre-processing, i.e. where the dataset is processed in such a manner that it produces less biased outcomes, before passing it to a model for training; 2) In-processing, i.e. where the model learns the original data distribution and shifts the data distribution to a fair distribution by adding some constraints during the training process; and 3) Post-processing, i.e. where the model’s outcome is changed in such a manner that it gives fair outcomes relative to protected attributes. The performance of these models or datasets can be measured by the evaluation metrics that reflect both the fairness and data utility. To ease up the work for training models and evaluating them, researchers has developed benchmarking tool that bring the training and evaluation in one framework. Recently, research on fair generative models has found a lot of spotlight and measuring the quality of the synthetic data is as crucial as evaluating fairness and data utlity.
@@
Figure 1: A High-level overview of FairX. An input dataset (possibly custom) is fed to the FairX data loading module followed by a debiasing module and an extensive evaluation module providing multi-faceted evaluations.
@@
The remainder of this paper is organised as follows. In Section 2 we discuss some background information that will help the reader understand the rest of the paper. We then present FairX in Section 3. Section 4 shows some fairness results obtained by FairX for a number of datasets and models. Details on how to use FairX are presented in Section 5. Finally, the paper looks ahead towards future improvements in Section 6.
@@ 2.3 Comparison of existing benchmarking tools
Over the years researchers have developed various fairness benchmarking tools which commonly include a dataset loader, different bias removal techniques and evaluation metrics. Fairlearn [2] by Microsoft is one such benchmarking tool. It has support for different algorithms for bias removal and measuring the fairness of a model. AIF360 [3] by IBM is another benchmarking tool. It supports a wide range of evaluation metrics (both for fairness and data utility) and bias-removal algorithms (in-processing, pre-processing and post-processing). Another example is Jurity [12]. It contains recommender system evaluations, and various fairness and data utility functions. More recently, in the area of generative models, there has been an increased interest in generating fair data in the image, tabular and medical domains [13, 14, 15, 1, 16, 17]. But the aforementioned benchmarking tools do not contain these models. Also, when evaluating models, other benchmarking tools, only measure the fairness and data utility of the models itself. But evaluation methods for generated data is needed. We need to verify the quality of the synthetic data. We also need to verify the authenticity of the synthetic data, to show the generative models are actually generating new content rather than just copying the data itself. FairX is bridging this gap. We add support for evaluating synthetic data and add generative models in our benchmarking tool. Table 1 shows the comparison among the models with FairX.
@@ 3 FairX
In this section we present FairX in detail. FairX is built on three primary modules, 1) the Data Loading Module, 2) the Bias-mitigating Techniques Module, and 3) the Evaluation Module. The main pipeline (shown in Figure 1) works as follows. Given a dataset, FairX will pre-processed it in a way that is compatible with the benchmarking model. Next the model will train itself using the dataset. After the training the evaluation module will give the results based on fairness, data utility and explain the outcome using explainability.
@@ 3.1 Data loading module
The BaseDataClass handles the internal processing of datasets and make it compatible with the bias-mitigating models that are present on our framework as well as making it easier to handle for other bias-mitigating models that are not present in this tool. This class contains different methods for handling different kinds of data extension (CSV, and others). We add three widely used tabular datasets (Adult-Income, COMPASS and Credit Card) and two image datasets (Colored MNIST and Celeba) in the benchmarking tool, and we plan to add more. Upon calling the BaseDataClass for a specific dataset, it downloads the dataset from a predefined URL. Then it processes it based on numerical and categorical features. The BaseDataClass provides methods to normalize the dataset and is equipped with functionality for various encodings (e.g. One-hot encoding, QuantileTransformer). It also has a dataset-splitting function to split the dataset for training and testing purpose. We also added functionality to prepare the dataset for explainability algorithms. Sample usage of datasets are described in section 5, Listing 1.
@@
Besides adding widely used benchmarking datasets for fair data research, we also provide the option to use custom dataset. By using the CustomDataClass, users can load their own dataset (CSV, TXT, etc.) and train the models. Users need to specify the sensitive attributes and target attributes while using the CustomDataClass. Pre-processing and other functionalities are also available in this class, like in the BaseDataClass. We present sample usage of CustomDataClass in Listing 4 of section 5.
@@ 3.2 Bias-mitigating techniques module
One of FairX’s main aims is to benchmark different bias-mitigation techniques on various datasets. Over the years, different techniques have been proposed, and we add models from these techniques to the tool. For the benchmarking process, we use the same hyper-parameters used in their respective works. We create a common format for all the bias-mitigation techniques to make it easy for the users. For example, each bias-mitigation technique has its own class, which has model.fit() function. This fit() function takes the dataset and processes it (if needed for the specific model). For the generative models (in-processing techniques), this function also generates synthetic data and saves it as a Pandas dataframe. Sample usage of models is described in section 5, Listing 2.
@@
We add the support for Correlation remover (CorrRemover in FairX) in the benchmarking. Correlation Remover removes the correlation between the sensitive attributes with other data features by using a linear transformation while keeping as much information as possible. It is also possible to control on how much correlation we want remove by using the remove_intensity parameter while the value $1.0$ will result maximum correlation removal while $0.0$ will do the opposite. We can access the pre-processing algorithm by using fairx.models.preprocessing.
@@
Most recent in-processing bias removal techniques are based on generative models. And the fairness benchmarking tools we mentioned in this work does not contain these models. One of our contribution of FairX is that, we add several fair generative models, such as, TabFairGAN [16], Decaf [17] and Fairdisco [1]. We can access the in-processing algorithm by using fairx.models.inprocessing module. After training, these models will generate and save the samples automatically.
@@
For the post-processing bias removal technique, We add Threshold Optimizer. This technique operates on a classifier and improve the output of its based on a fairness constraint. In this case, we use demographic_parity as a fairness constraint to improve the outcome of the classifier. For using the post-processing algorithm, we can use fairx.models.postprocessing module.
@@
Adult-Income, Compass, Credit-card
@@ 4 Results and discussion
Table 3 and 4 shows the performance of the bias removal algorithm for the Adult-Income dataset and Compass dataset respectively. We run experiment using different Protected attributes [^1] For the sake of brevity, we could not include additional results using other datasets—we refer the reader to the FairX repository for these results. Some metrics like precision, recall, fairness through unawareness (FTU), and plots like fairness-accuracy trade-offs were similarly omitted.. Besides, fairness and data utility, we add synthetic data evaluation for the output of TabFairGAN, and Decaf.
Figure 2: PCA and t-SNE plots of the original data and Synthetic data generated by TabFairGAN. Here each dot represents a record, if the generative model learns the original data distribution then the dots should overlap with each other. Dataset: Adult-Income, Protective attribute: ‘sex’.
From the table, we see for the generative fair models, TabFairGAN is performing well comparing with the Decaf in both datasets with both protected attributes. The $\alpha-$precision, $\beta-$recall scores of TabFairGAN is better than Decaf, this represents the synthetic data quality of TabFairGAN is superior than Decaf. On the other hand, TabFairGAN perform poorly in the fairness evaluation for the ‘race’ protected attribute of the Adult-Income dataset. Pre-processing technique Correlation Remover and in-processing technique FairDisco also perform well in terms of fairness and data utility.
@@
5 Usage
@@
6 Conclusion and future work
Massive amounts of data are being produced everyday. Unfortunately, much of this data contains human or machine biases. Furthermore, the usage of recommendation system has increased with advancements in artificial intelligence. But if we use biased data to train a recommendation system, there is a high chance that the recommendation system will yield unfair decision towards some demographics. To mitigate this issue, researchers have developed various measure to remove the bias from the dataset or models; or to train the model in such a way that the model produces bias-free data. To help in this process, benchmarking tools equipped with different bias-removal techniques and evaluation metrics were developed over the years. But these benchmarking tools commonly lack the option to evaluate generative models or to train them. We therefore presented FairX, an open-source, modular, fairness benchmarking tool. FairX comes with a data-loader, supports model training, and has an evaluation module. FairX provides support for training fair generative models and for evaluating the synthetic data created by them. FairX also contains various fairness evaluation metrics, data utility evaluation metrics and different plotting techniques to help users to evaluate models and visualize outcomes. FairX comes with support for explainability analysis of a prediction using the dataset (both original and synthetic) and shows feature importance. We believe FairX will help the researchers and mitigate the gap of not having fair generative models and way of evaluating synthetic data.
Acknowledgements.

### v4

# Accepted at AEQUITAS 2024: Workshop on Fairness and Bias in AI | co-located with ECAI 2024, Santiago de Compostela, Spain

@@ Abstract
We present FairX, an open-source Python-based benchmarking tool designed for the comprehensive analysis of models under the umbrella of fairness, utility, and eXplainability (XAI). FairX enables users to train benchmarking bias-mitigation models and evaluate their fairness using a wide array of fairness metrics, data utility metrics, and generate explanations for model predictions, all within a unified framework. Existing benchmarking tools do not have the way to evaluate synthetic data generated from fair generative models, also they do not have the support for training fair generative models either. In FairX, we add fair generative models in the collection of our fair-model library (pre-processing, in-processing, post-processing) and evaluation metrics for evaluating the quality of synthetic fair data. This version of FairX supports both tabular and image datasets. It also allows users to provide their own custom datasets. The open-source FairX benchmarking package is publicly available at https://github.com/fahim-sikder/FairX.
@@ 1 Introduction
With the rapid development of artificial intelligence-based systems to aid us in our daily lives, it is important for these systems to give outcomes that is acceptable for all users, including—but not limited to—from demographic perspective. Troublingly, as the available data is filled with human or machine bias, models trained with these dataset often gives unfair outcome towards some demographic [1]. It is therefore critical to mitigate bias in the dataset and model. Over the years, researchers have used different techniques to achieve this [2, 3]. These techniques can be roughly grouped into three families: 1) Pre-processing, i.e. where the dataset is processed in such a manner that it produces less biased outcomes, before passing it to a model for training; 2) In-processing, i.e. where the model learns the original data distribution and shifts the data distribution to a fair distribution by adding some constraints during the training process; and 3) Post-processing, i.e. where the model’s outcome is changed in such a manner that it gives fair outcomes relative to protected attributes. The performance of these models or datasets can be measured by the evaluation metrics that reflect both the fairness and data utility. To ease up the work for training models and evaluating them, researchers has developed benchmarking tool that bring the training and evaluation in one framework. Recently, research on fair generative models has found a lot of spotlight and measuring the quality of the synthetic data is as crucial as evaluating fairness and data utlity.
@@
Figure 1: A High-level overview of FairX. An input dataset (possibly custom) is fed to the FairX data loading module followed by a bias-mitigation module and an extensive evaluation module providing multi-faceted evaluations.
@@
The remainder of this paper is organised as follows. In Section 2 we discuss some background information that will help the reader understand the rest of the paper. We then present FairX in Section 3. Section 4 shows some fairness results obtained by FairX for a number of datasets and models. Finally, the paper looks ahead towards future improvements in Section 5.
@@ 2.3 Comparison of existing benchmarking tools
Over the years researchers have developed various fairness benchmarking tools which commonly include a dataset loader, different bias mitigation techniques and evaluation metrics. Fairlearn [4] by Microsoft is one such benchmarking tool. It has support for different algorithms for bias mitigation and measuring the fairness of a model. AIF360 [5] by IBM is another benchmarking tool. It supports a wide range of evaluation metrics (both for fairness and data utility) and bias-removal algorithms (in-processing, pre-processing and post-processing). Another example is Jurity [14]. It contains recommender system evaluations, and various fairness and data utility functions. AEQUITAS [15], FairBench [16] generate fairnes report and REVISE [17] is a tool to detect and mitigate bias in the image dataset. More recently, in the area of generative models, there has been an increased interest in generating fair data in the image, tabular and medical domains [18, 19, 20, 1, 21, 22]. But the aforementioned benchmarking tools do not contain these models. Also, when evaluating models, other benchmarking tools, only measure the fairness and data utility of the models itself. But evaluation methods for generated data is needed. We need to verify the quality of the synthetic data. We also need to verify the authenticity of the synthetic data, to show the generative models are actually generating new content rather than just copying the data itself. FairX is bridging this gap. We add support for evaluating synthetic data and add generative models in our benchmarking tool. Table 1 shows the comparison among the models with FairX.
@@ 3 FairX
In this section we present FairX in detail. FairX is built on three primary modules, 1) the Data Loading Module, 2) the Bias-mitigating Techniques Module, and 3) the Evaluation Module. The main pipeline (shown in Figure 1) works as follows. Given a dataset, FairX will pre-process it in a way that is compatible with the benchmarking model. Next the model will train itself using the dataset. After the training the evaluation module will give the results based on fairness, data utility and explain the outcome using explainability.
@@ 3.1 Data loading module
The BaseDataClass handles the internal processing of datasets and make it compatible with the bias-mitigating models that are present on our framework as well as making it easier to handle for other bias-mitigating models that are not present in this tool. This class contains different methods for handling different kinds of data extension (CSV, and others). We add three widely used tabular datasets (Adult-Income, COMPAS and Credit Card) and two image datasets (Colored MNIST and CelebA) in the benchmarking tool, and we plan to add more. The BaseDataClass process datasets based on numerical and categorical features. It also provides methods to normalize the dataset and is equipped with functionality for various encodings (e.g. One-hot encoding, QuantileTransformer). It also has a dataset-splitting function to split the dataset for training and testing purpose. We also add functionality to prepare the dataset for explainability algorithms. Sample usage of datasets are described in Appendix Section A, Listing 1.
@@
Besides adding widely used benchmarking datasets for fair data research, we also provide the option to use custom dataset. By using the CustomDataClass, users can load their own dataset (CSV, TXT, etc.) and train the models. Users need to specify the sensitive attributes and target attributes while using the CustomDataClass. Pre-processing and other functionalities are also available in this class, like in the BaseDataClass. We present sample usage of CustomDataClass in Listing 4 of Appendix Section A.
@@ 3.2 Bias-mitigating techniques module
One of FairX’s main aims is to benchmark different bias-mitigation techniques on various datasets. Over the years, different techniques have been proposed, and we add models from these techniques to the tool. For the benchmarking process, we use the same hyper-parameters used in their respective works. We create a common format for all the bias-mitigation techniques to make it easy for the users. For example, each bias-mitigation technique has its own class, which has model.fit() function. This fit() function takes the dataset and processes it (if needed for the specific model). For the generative models (in-processing techniques), this function also generates synthetic data and saves it as a Pandas dataframe. Sample usage of models is described in Appendix Section A, Listing 2.
@@
We add the support for Correlation remover [4] (CorrRemover in FairX) in the benchmarking. Correlation Remover removes the correlation between the sensitive attributes with other data features by using a linear transformation while keeping as much information as possible. It is also possible to control on how much correlation we want remove by using the remove_intensity parameter while the value $1.0$ will result maximum correlation removal while $0.0$ will do the opposite. We can access the pre-processing algorithm by using fairx.models.preprocessing.
@@
Most recent in-processing bias mitigation techniques are based on generative models. And the fairness benchmarking tools we mentioned in this work does not contain these models. One of our contribution of FairX is that, we add several fair generative models, such as, TabFairGAN [21], Decaf [22] and Fairdisco [1]. We can access the in-processing algorithm by using fairx.models.inprocessing module. After training, these models will generate and save the samples automatically.
@@
For the post-processing bias mitigation technique, We add Threshold Optimizer [4]. This technique operates on a classifier and improve the output of its based on a fairness constraint. In this case, we use demographic_parity as a fairness constraint to improve the outcome of the classifier as presented in [4]. For using the post-processing algorithm, we can use fairx.models.postprocessing module.
@@
Adult-Income, Compas, Credit-card
@@ 4 Results and discussion
Table 3 and 4 shows the performance of the bias mitigation algorithms for the Adult-Income dataset and Compas dataset respectively. We run experiment using different Protected attributes  [^1] For the sake of brevity, we could not include additional results using other datasets—we refer the reader to the FairX repository for these results. Some metrics like precision, recall, fairness through unawareness (FTU), and plots like fairness-accuracy trade-offs were similarly omitted.. Besides, fairness and data utility, we add synthetic data evaluation for the output of TabFairGAN  [^2] https://github.com/amirarsalan90/TabFairGAN, and Decaf  [^3] https://github.com/vanderschaarlab/synthcity.
Figure 2: PCA and t-SNE plots of the original data and Synthetic data generated by TabFairGAN. Here each dot represents a record, if the generative model learns the original data distribution then the dots should overlap with each other. Dataset: ‘Adult-Income’, Protective attribute: ‘sex’.
From the table, we see for the generative fair models, TabFairGAN is performing well comparing with the Decaf in both datasets with both protected attributes. The $\alpha-$precision, $\beta-$recall scores of TabFairGAN is better than Decaf, this represents the synthetic data quality of TabFairGAN is superior than Decaf. On the other hand, TabFairGAN perform poorly in the fairness evaluation for the ‘race’ protected attribute of the Adult-Income dataset. Whereas In-processing technique FairDisco  [^4] https://github.com/SoftWiser-group/FairDisCo performs well in terms of fairness and data utility.
@@
5 Conclusion and future work
Massive amounts of data are being produced everyday. Unfortunately, much of this data contains human or machine biases. Furthermore, the usage of recommendation system has increased with advancements in artificial intelligence. But if we use biased data to train a recommendation system, there is a high chance that the recommendation system will yield unfair decision towards some demographics. To mitigate this issue, researchers have developed various measure to mitigate the bias from the dataset, or to train the model in such a way that the model produces bias-free data. To help in this process, benchmarking tools equipped with different bias-mitigation techniques and evaluation metrics were developed over the years. But these benchmarking tools commonly lack the option to evaluate generative models or to train them. We therefore presented FairX, an open-source, modular, fairness benchmarking tool. FairX comes with a data-loader, supports model training, and has an evaluation module. FairX provides support for training fair generative models and for evaluating the synthetic data created by them. FairX also contains various fairness evaluation metrics, data utility evaluation metrics and different plotting techniques to help users to evaluate models and visualize outcomes. FairX comes with support for explainability analysis of a prediction using the dataset (both original and synthetic) and shows feature importance. We believe FairX will help the researchers and mitigate the gap of not having fair generative models and way of evaluating synthetic data.
Acknowledgements.
Appendix A Detailed Usage